You are trying to find out how high you have
to pitch a water balloon in order for it to burst
when it hits the ground. You discover that
the balloon bursts when you have pitched it
to a height of 14 m.
With what velocity did the balloon hit
the ground? The acceleration of gravity is
9.8 m/s
2
.
Answer in units of m/s
Answer #1
Vf^2 = Vo^2 + 2g*d,
Vf^2 = 0 + 19.6*14 = 274.4,
Vf = 16.6m/s.
answered by: Lorena